### v3

# Hardware-Aware Parallel Prompt Decoding for Memory-Efficient Acceleration of LLM Inference Accepted at EMNLP 2025

Abstract
The auto-regressive decoding of Large Language Models (LLMs) results in significant overheads in their hardware performance. While recent research has explored various speculative decoding techniques for multi-token generation, these methods introduce high memory costs from the additional weights and KV cache of separate draft models, limiting efficiency in edge and long-context scenarios. To overcome these limitations in edge-scale LLMs, we propose a novel parallel prompt decoding that requires only runtime memory overhead by employing a unified single model for both speculation and verification. Inspired by the human natural language generation process, PPD approximates outputs generated at future timesteps in parallel by using multiple prompt tokens. Furthermore, we present a hardware-aware two-stage tree pruning algorithm that adaptively optimizes this decoding scheme to fully leverage the computational capacities on different GPUs. Through extensive experiments across LLMs ranging from MobileLlama to Vicuna-13B on a wide range of benchmarks, our approach demonstrates up to 2.49 speedup. Moreover, our parallel prompt decoding can serve as an orthogonal optimization for synergistic integration with existing speculative decoding, showing up to further speed improvement. To support future development, we have included our code implementation with this submission.
[…]
Hardware-Aware Parallel Prompt Decoding for
Memory-Efficient Acceleration of LLM Inference
††thanks: Accepted at EMNLP 2025
Hao (Mark) Chen1 Wayne Luk1 Ka Fai Cedric Yiu2 Rui Li3 Konstantin Mishchenko3 Stylianos I. Venieris3 Hongxiang Fan1,3 1Imperial College London, UK 2Hong Kong Polytechnic University, Hong Kong 3Samsung AI Center, Cambridge, UK {hc1620,w.luk}@ic.ac.uk {rui.li,s.venieris}@samsung.com konsta.mish@gmail.com cedric.yiu@polyu.edu.hk hongxiangfan@ieee.org
1 Introduction
The recent advances in large language models (LLMs) are increasingly shaping and influencing a wide range of AI applications. However, autoregressive generation, the de facto approach employed in LLM inference, suffers from inadequate hardware performance due to its inherent sequential nature (Stern et al., 2018). Speculative decoding (Leviathan et al., 2023; Chen et al., 2023; Kim et al., 2024), an emerging acceleration technique, employs a guess-and-verify framework for LLM inference, where a smaller draft model first predicts multiple tokens sequentially and then the original LLM verifies them in parallel. Despite its potential, speculative decoding is constrained by high training costs and memory overhead, as shown in Figure 1. Particularly, this paper reveals that the extra runtime memory required for draft model storage and the KV cache—which scales linearly with sequence length—limits its practicality in edge devices, mobile environments, and long-context scenarios. Therefore, given the growing need for user privacy and personalization, this paper focuses on accelerating edge-scale LLMs via a novel more memory- and cost-efficient acceleration solution.
Recent efforts have explored the possibility of generating multiple tokens in parallel without relying on a separate transformer draft model (Santilli et al., 2023). Approaches such as inserting additional decoding heads (Cai et al., 2024) and retrieving frequently used tokens (He et al., 2023) are employed to enhance performance. However, these methods either make strong conditional independence assumptions for tokens generated within a single step (Cai et al., 2024) or rely on unparameterized trained tokens that do not scale effectively (Lin et al., 2024). Therefore, they often suffer from low acceptance rates during inference.
To alleviate the complexity and overhead associated with the use of draft models while maintaining a high acceptance rate, we propose Parallel Prompt Decoding (PPD), a novel architecture-agnostic and memory-efficient framework that adopts prompt tuning for non-autoregressive LLM inference. Inspired by the human natural language generation process where continuous words like common expressions and phrases are produced simultaneously, PPD introduces the use of ensemble prompt tokens, the meticulously trained embeddings, for multi-token prediction. Specifically, these trained tokens are appended to the original input sequence in parallel, enabling the concurrent generation of multiple output tokens in a single forward pass. The key intuition of PPD lies in the observation that if trained properly, prompt tokens appended to the input can approximate tokens generated at future timesteps, thereby partially recovering the missing conditional dependency information for multi-token generation. By strategically positioning trained prompt tokens, PPD achieves up to a 28% higher acceptance rate when predicting long-range tokens. Moreover, to facilitate the optimized implementation of PPD across different hardware platforms, we propose a hardware-aware two-stage tree pruning technique that adaptively refines the prompt structure during runtime based on the computational resources available on the specific hardware. As shown in Figure 1(a), PPD achieves a comparable speedup to the state-of-the-art speculative decoding approaches with negligible memory overhead and reduced training cost.
To demonstrate the effectiveness of our approach, we evaluate PPD on MobileLLaMA (Chu et al., 2023), Vicuna-7b and Vicuna-13b (Chiang et al., 2023). Running on a single GPU using the A100-40GB and RTX 4090, our method achieves a speedup ratio for inference from 2.12 to 2.49 across a diverse range of popular datasets including MT-Bench, HumanEval, and GSM8K. Our experiments demonstrate that PPD not only achieves comparable throughput to the state-of-the-art speculative decoding method, but it also manages this with significantly fewer trainable parameters—specifically, 0.0002% of trainable parameters—and incurs only a minimal memory overhead (0.0004%), showcasing that PPD is remarkably cost- and memory-efficient. The training of prompt tokens can be completed in 16 hours using one A100 GPU, 8 hours using four GeForce RTX 3090 GPUs, compared to the 1-2 days on four A100 GPUs required for Eagle (Li et al., 2024b). Furthermore, since PPD does not require the modification of the original LLM or the addition of extra networks, it is highly adaptable and orthogonal to other decoding techniques. For instance, it can be effectively combined with a draft model to further reduce inference latency.
Our contributions are summarized as follows:
A novel Parallel Prompt Decoding (PPD) that adopts cost-effective prompt tokens for non-autoregressive LLM inference, achieving a high acceptance rate for long-distance token prediction with preserved output quality.
A hardware-aware two-stage tree pruning technique that adaptively optimizes the prompt structure of PPD at runtime based on the available compute and memory resources, facilitating its efficient deployment on various hardware platforms.
An open-source implementation of PPD, accompanied by comprehensive evaluations on various models and benchmarks. Our experiments demonstrate that PPD achieves significant speed improvements with negligible memory overhead and reduced training cost.
2 Background and Related Work
To enhance the inference speed of LLM, various approaches adopt an iterative guess-and-verify strategy to enable multi-token generation. In the guessing phase, potential future tokens are proposed at a faster speed than in traditional autoregressive implementations. Subsequently, a parallelized verification process assesses which guessed tokens should be accepted. Depending on how tokens are generated during the guess stage, these approaches can generally be categorized as i) speculative decoding and ii) parallel decoding.
2.1 Speculative Decoding
The guessing phase of speculative decoding adopts a lightweight draft model to generate multiple tokens at an increased speed (Kim et al., 2024). During the verification stage, the original LLM determines the acceptance of the guessed tokens. It is worth noting that both draft and original models still follow the auto-regressive inference scheme.
Building on the speculative decoding scheme, various studies have been conducted to further optimize its inference speed. To improve the accuracy of the draft model, Eagle (Li et al., 2024b) incorporates the hidden features into the draft model’s forward pass. Recently, Eagle-2 (Li et al., 2024a) enhances their approach using a context-aware dynamic tree construction. However, both Eagle and Eagle-2 utilize a separate draft model for multi-token generation, diverging fundamentally from our prompt decoding approach. Moreover, their dynamic tree construction scheme is an orthogonal technique to our two-stage tree pruning method. SpecInfer (Miao et al., 2024) adopts a tree-based speculative inference and verification scheme, improving the diversity of speculation candidates. Sequoia (Chen et al., 2024) optimizes the sparse tree structure of speculative decoding by considering the capability of the underlying hardware platforms. Our tree pruning algorithm differs from Sequoia by accounting for two types of tokens in the tree: prompt tokens and guess tokens, whereas Sequoia only considers guess tokens.
2.2 Parallel Decoding
To overcome the inherent limitations of autoregressive inference and the memory overhead associated with using a separate draft model, several attempts have been made to integrate both guessing and verification using one unified model. Medusa111We categorize Medusa as parallel decoding because it only adopts LM heads instead of separate models. (Cai et al., 2024) introduces language model (LM) heads at the final layer of the original LLM, facilitating the generation of multiple tokens in a single forward pass. It also utilizes tree attention masks in its verification process to increase speed even further. To enhance token drafting with retrieval-augmented generation (Karpukhin et al., 2020), Rest (He et al., 2023) introduce retrieval-based decoding tailored for specific scenarios. Inspired by Jacobi decoding (Santilli et al., 2023) that adopts multiple special tokens to accelerate machine translation, Lookahead Decoding (Fu et al., 2024) improves upon this method by generating parallel n-grams and employing a caching memory pool. To capture more information while using multiple special tokens at distinct positions, PaSS (Monea et al., 2023) and BiTA Lin et al. (2024) train additional token embeddings for parallel decoding. Compared with these parallel decoding methods, our proposed PPD introduces two key innovations that distinguish it: 1) PPD trains the embeddings of parameterized ensemble prompt tokens, 2) it utilizes a hardware-aware two-stage tree pruning algorithm for designing a sparse tree tailored to each hardware platform.
3 Parallel Prompt Decoding (PPD)
The primary advantage of PPD lies in training embeddings for prompt tokens rather than developing a separate model. Our method integrates three substeps into a single decoding step, following the guess-and-verify strategy: (1) candidate generation, where multiple candidate continuations222A candidate token, also referred to as a ”guess token”, is a draft token generated from a prompt token. are predicted by strategically inserting the prompt tokens into the input sequence. We adopt tree attention (Miao et al., 2024) to merge the processing of multiple candidates into a single forward pass; (2) candidate verification, where two verification schemes, exact matching (Fu et al., 2024) and typical acceptance (Cai et al., 2024), are implemented; (3) candidate acceptance, where validated candidates are integrated into the input and KV cache is updated accordingly. Figure 2 presents the inference scheme of combining generation and verification steps in a single forward pass.
3.1 Prompt Tokens
The prompt tokens are the key component of PPD to realize multi-token generation. Initially introduced by Lester et al. (2021) to adapt LLMs for specific tasks, prompt tokens are typically prepended to the input, with outputs generated in an autoregressive manner. In this work, we propose a novel approach of utilizing prompt tokens by strategically positioning them at locations where tokens are anticipated to be generated in parallel.
In the standard decoding process, the probability of predicting the next token is expressed as the conditional probability , where is the input prompt, are the tokens generated so far, and is the next token to be predicted. For conventional parallel decoding techniques (Stern et al., 2018; Cai et al., 2024) that presume complete conditional independence among tokens decoded in a single step, the exact conditional probability is approximated by , where indicates the token distance.333The token distance is the number of tokens between the last accepted token and the predicted token. However, we observe that as increases, the gap between the actual probability and its approximation expands due to the absence of relevant conditional dependencies. We argue that prompt tokens can bridge this gap by more accurately modeling the conditional probability as , where is the prompt token with token distance . Through this forward pass in the decoder layers, these causally linked prompt tokens facilitate the flow of information along the sequence of speculative tokens, restoring the conditional probability. We demonstrate the effectiveness of this approach in Section 5.2.
3.2 Ensemble Prompt Tokens
Inspired by prompt ensembling (Lester et al., 2021), which uses multiple prompts to generate diverse responses and aggregates these to derive a single answer, we introduce the concept of ensemble prompt token (EPT). This abstraction allows us to decouple each prompt token from the fixed embedding dimension. For every prompt token, there exist multiple corresponding EPTs, each with its distinct embedding. We modify the attention mask to ensure that each EPT only depends on the corresponding EPTs from preceding prompt tokens. This selective visibility is maintained for both training and inference, where the guess token for each prompt token is determined by averaging the logits of its EPTs. The use of EPTs not only enables direct and flexible control over the trainable parameters, but also leads to increased prediction accuracy. The probability is approximated as , where denotes the EPT at a token distance of . Further details can be found in Appendix F.
3.3 Training
During training, only the embeddings of prompt tokens are changed, with the parameters of the original LLM remaining frozen. We adopt the following two training techniques:
Random Insertion of Prompt Tokens: Randomly inserting prompt tokens throughout the input sequence reduces contextual bias from appending them only at the end. This approach broadens the predictive capacity of prompt tokens beyond a limited vocabulary such as <eos> and punctuation.
Knowledge Distillation: To align the predictive behavior of prompt tokens with the original LLM, we employ knowledge distillation. Instead of using hard labels, prompt tokens are trained against the logits produced by the original LLM. Following Medusa (Cai et al., 2024), the loss function is formulated as:
where is the KL divergence, is the predicted distribution of the prompt token, is the corresponding distribution from the original LLM, and is the decay ratio.
4 Sparse Tree Pruning
4.1 Customized Sparse Tree Attention
Tree attention, introduced by SpecInfer (Miao et al., 2024), increases the expected acceptance rate by considering the top-k candidates from a single decoding step. In their approach, the input is structured as a tree, where each level of the tree corresponds to a specific output position. An attention mask is applied to the tree-structured input, allowing the model to process multiple candidates efficiently without increasing the batch size.
To improve the efficiency and performance of LLM inference, this paper proposes a novel sparse tree customized for PPD, which prioritizes candidates in the tree structure with higher prediction accuracy. A key difference from previous works (Cai et al., 2024; Chen et al., 2024) is the appending of a sequence of prompt tokens to each guess token. The length of the prompt token sequence decides the maximum depth of the speculative tree at the next decoding step. To further hide the latency introduced by the extra prompt tokens, we propose a novel tree pruning algorithm (Section 4.2) that optimizes the number of prompt tokens.
4.2 Two-Stage Tree Pruning Algorithm
As depicted in Figure 3, our tree pruning algorithm consists of two stages: an offline static tree pruning phase and an online hardware-aware tree optimization phase. These two stages are applied subsequently to reduce the amount of computation involved in PPD multi-token generation.
Static Tree Pruning. The first stage, static tree pruning, is applied offline prior to runtime deployment. The goal is to reduce the number of prompt tokens in the tree to achieve the desired tree size. As shown on the left side of Figure 3, the tree pruning process consists of three key steps:
- 1.
2.
Prompt Tokens Appending : Attaching the maximum allowable prompt tokens to each candidate token from the first step.
3.
Greedy Prompt Token Removal: Removing a prompt token greedily to maximize expected amortized acceptance lengths, continuing until the desired prompt token budget is reached.
Each guess token in the tree is appended with a sequence of prompt tokens, with each prompt token corresponding to a unique output position. The length of this sequence determines the tree’s maximum depth at the next decoding step. Thus, removing a prompt token at a guess token reduces the maximum tree depth at the next decoding step if this guess token is accepted in the current step. Let represent the acceptance probability of guess token , and denote the expected acceptance length with prompt tokens before removal. The decrease in expected acceptance length, , due to removing a prompt token at is given by . More details are discussed in Appendix B.
Hardware-Awareness Tree Optimization. Given that hardware platforms differ in terms of memory and compute resources, we propose a hardware-aware tree optimization to maximize the performance of PPD. As shown on the right of Figure 3, this optimization adjusts the tree size budget based on the performance characteristics of the target hardware.
To achieve this, we define two key functions: 1. Acceptance length (hardware-independent) and 2. Forward pass latency (hardware-dependent). The speedup ratio, , is estimated using a validation dataset, with evaluated once and tested on different hardware platforms. We then choose the tree size budget that maximizes based on the measured runtime latency on the specific hardware platform. To eliminate runtime overhead, hardware latency profiling is conducted during idle periods.
5 Experiments
Models and testbeds. We conducted all the experiments using MobileLLaMA-1.4B (Chu et al., 2023), Vicuna-7B and Vicuna-13B (Chiang et al., 2023). We used 3 prompt tokens and 1 EPT per prompt token for all inference experiments. The inference throughputs of the models are evaluated on a single NVIDIA A100 GPU with 40GB of memory and a GeForce RTX 4090 using a batch size of 1 and FP16 precision. Further details about the experimental setup can be found in Appendix H.
Training. We froze all trainable parameters of the original LLM. Prompt token embeddings were trained using distillation logits generated from the ShareGPT dataset (ShareGPT, 2023), with a maximum context length of 1024, a cosine learning rate scheduler starting at 0.01, and no warmup. Prompt token embeddings are initialized with normal text token embeddings. For each model, the same set of prompt tokens is used across experiments to demonstrate its generalizability.
Datasets. We assess the throughput performance of PPD across various tasks and datasets. Specifically, we evaluated PPD using the MT-Bench dataset (Zheng et al., 2023), which contains multi-turn questions with a range of topics, in both non-greedy (temperature follows the default configuration) and greedy settings (temperature=0). We used the GSM8K (Cobbe et al., 2021) and HumanEval (Chen et al., 2021) datasets only in the greedy setting. The GSM8K dataset consists of grade school math problems and we used the first 500 questions of the test split for our evaluations. HumanEval includes coding tasks, for which we set a maximum new token limit of 512 to control the length of the generated sequences. We used the Alpaca (Li et al., 2023) dataset as the validation dataset to produce the latencies and acceptance lengths used for sparse tree pruning.
5.1 Speedup Comparison with Parallel Decoding Methods
We compare the speedup ratios of PPD with state-of-the-art parallel decoding methods on MT-Bench in non-greedy settings in Figure 4. PPD achieves speedups up to 13.8% higher than Medusa and between 2 times and 3 times higher than other parallel decoding methods. We examine the factors contributing to the enhanced speedup ratios and other performance metrics, as presented in Table 1. The reasons for the increase in speedup ratios are two-fold. Firstly, PPD produces candidate tokens with a higher acceptance rate than Medusa when utilizing a sparse tree of the same size. Notably, PPD continues to achieve a comparable or slightly better acceptance rate even when employing a much smaller sparse tree – ranging from one-third to half the size. Secondly, PPD benefits from lower forward pass latency due to its ability to use smaller sparse tree sizes and hence shorter input lengths. PPD also eliminates the computational overhead associated with separate decoding heads. PPD maintains the same output quality, achieving about the same score on MT-Bench while using significantly fewer trainable parameters.
| Model | Method | (s) | Quality | (%) | ||||
| M | Vanilla | 50.2 | 1.00 | 0.020 | - | NA | NA | 1 |
| PPD | 108.7 | 2.43 | 0.022 | Same | 4.50 | (10,84,89) | (40,285,285) | |
| V-7B | Vanilla | 39.2 | 1.00 | 0.026 | 5.99 | NA | NA | 1 |
| Medusa | 82.0 | 2.51 | 0.0307 | 5.98 | 8.07 | 63 | 63 | |
| PPD | 88.0 | 2.54 | 0.029 | 5.93 | 1.82 | (10,33,34) | (40,105,105) | |
| V-13B | Vanilla | 30.4 | 1.00 | 0.0330 | 6.38 | NA | NA | 1 |
| Medusa | 63.4 | 2.59 | 0.0408 | - | 5.52 | 63 | 63 | |
| PPD | 66.1 | 2.44 | 0.0379 | 6.32 | 7.87 | (10,20,20) | (40,60,60) |
Figure 5 displays the throughput of PPD on MT-Bench, HumanEval, and GSM8K with temperature equal to 0. PPD achieves consistent walltime speedup ratios from 2.12 to 2.49 on different GPUs, which demonstrates that prompt tokens generalize well on different tasks. In general, PPD performs better in coding and math reasoning tasks, achieving speedups between 2.21 and 2.49. This can be attributed to the fact that both code and math equations often contain fixed patterns and repetitive symbols, which narrows the range of plausible candidates and simplifies the prediction. We also found that with typical acceptance, the speedup increases with temperature. Another notable trend is that smaller models, such as Vicuna-7B, generally achieve more significant speedup ratios as compared to larger models, like Vicuna-13B. PPD aims to generate more tokens per step, which comes with increased computational demands. For larger models that already require substantial computational resources, it is necessary to limit the size of the sparse tree to avoid exceeding the GPU’s utilization cap and causing increased latency. As a result, the number of tokens accepted per step is reduced, leading to lower speedups. However, this can be amortized when using more powerful GPUs than the NVIDIA A100 and the RTX 4090, such as NVIDIA H100.
5.2 Long-range Token Prediction
For a specific sparse tree, the accumulative accuracy provides a theoretical upper bound for the number of generated tokens per step and the maximum possible speedup ratio. Hence, maximizing accumulative accuracy is crucial for the effectiveness of PPD. Figure 6 demonstrates the accumulative accuracy of the tokens predicted at various positions. We summarize the following three key insights from the results.
PPD excels at predicting more distant tokens. As depicted in Figure 6(a), PPD consistently outperforms Medusa in accuracy across all token positions. The accuracy gap between PPD and Medusa widens with the increased token distance (e.g., the top-10 accuracy difference is 0.03 for the ‘next next’ word versus 0.12 for the ‘next next next next’ word). This improvement can be attributed to PPD’s ability to partially recover conditional dependency information through causally connected prompt tokens.
PPD performs well at generating a broader array of plausible token candidates. For example, in predicting the token at a token distance of 3, the top-10 candidates exhibit an accuracy improvement of 0.1 over Medusa, compared to only 0.02 for the top-1 candidate. This demonstrates the value of using tree attention and the largest viable tree size during inference, as multiple candidate continuations further boost accuracy improvement.
Multiple EPTs per prompt token and larger model sizes yield modest improvements in prediction accuracy. Figure 6(b) shows that using 100 EPTs per prompt token leads to accuracy improvement, ranging from 0.018 to 0.045. Figure 6(c) displays that PPD with Vicuna-13B outperforms Vicuna-7B with an accuracy gain of 0.0110.038. This increase is due to Vicuna-13B’s greater embedding dimensions and deeper layers, which enhance the expressive power of prompt tokens. However, these gains are modest and can be offset by the increased computational burden of larger models.
5.3 Memory and Training Efficiency
Memory efficiency. As shown in Figure 7(a), we compare the memory overhead of PPD with the leading parallel decoding (Medusa) and speculative decoding approaches (Eagle). The memory overhead of PPD is negligible, whereas Eagle incurs up to 5.8 overhead and Medusa up to 1.6. This efficiency stems from: (1) PPD’s parameter efficiency—token embeddings are much smaller than decoding heads and draft models, which scale with vocabulary size; and (2) using a single KV cache, unlike methods like Eagle that require separate caches for the draft and target models.
| Method | Training Time |
| PPD (Ours) | 0.52 hours |
| Medusa | 1.24 hours |
| Eagle | 1-2 days |
Training efficiency. Table 2 compares the training times of PPD with parallel and speculative decoding methods. PPD is trained until its evaluation accuracy of top-10 candidates surpasses that of Medusa on Alpaca Eval. Notably, PPD surpasses Medusa in evaluation accuracy while training in less than half the time, demonstrating its great potential to reduce training cost.
5.4 Ablation Study
Tree Attention. As illustrated in Figure 7(b), tree attention boosts the speedup ratio of PPD by an additional 32%, indicating that PPD generates accurate top-k predictions. Even without the use of tree attention, PPD still outperforms all other parallel decoding methods, achieving up to a 14% higher speedup ratio, demonstrating the effectiveness.
Sparse Tree Pruning Algorithm. Figure 8(a) shows that the pruned sparse trees consistently achieve longer acceptance lengths compared to naive and random ones across varying sizes. The acceptance length for pruned sparse trees shows a steady increase as the tree size extends, suggesting its good scalability. The convergence of pruned and naive sparse trees at larger sizes suggests a structural similarity emerging from constraints in tree depth and tree node count.
Hardware-aware Tree Size. Figure 8(b) presents the theoretical speedup across different GPUs. Figure 8(c) validates that the optimal sparse tree size, derived from theoretical speedup models, indeed results in the greatest actual speedup observed.
PPD + Speculative Decoding. As an orthogonal optimization in accelerating LLMs, PPD can be easily integrated with speculative decoding (Kim et al., 2024). To demonstrate this, we applied PPD to Vicuna-68M (Yang et al., 2024) and used it as the draft model for Vicuna-7B. This combination resulted in a speedup of up to 1.22 for speculative decoding on Vicuna-7B compared to using speculative decoding alone.
6 Conclusion
We introduced PPD, a memory-efficient and cost-effective parallel decoding method that incorporates a hardware-aware online tree optimization. Utilizing specially trained prompt tokens to predict long-range tokens accurately, PPD achieves a speedup of up to 2.49 in inference while employing negligible additional trainable parameters and memory overhead. We showcased that PPD offers a novel perspective on the capabilities of parallel decoding. Importantly, it could be synergized with other speculative or parallel decoding techniques to expedite inference even further.
Limitations
Despite its efficiency, we have identified the following limitations of PPD:
1.
GPU compute resource constraint. Since PPD trades additional compute resources for increased throughput, its effectiveness depends on the availability of idle GPU compute resources. On a GPU with limited compute resources, the speedup ratios achieved by PPD are expected to decrease.
2.
Extended input length. The improvement in acceptance length with PPD is not as significant as the gain in prediction accuracy compared to Medusa. This is because PPD must reserve a substantial portion of the input for prompt tokens, which limits the size of the sparse tree that can be used.
Appendix A Appendix
Appendix B Detailed Tree Construction Algorithm
We follow the same optimal sparse tree construction approach as in Medusa (Cai et al., 2024).
Definition B.1.
Let be the maximal number of prompt tokens per tree node. The sparse tree can exist in states, each represented by corresponding to state , where . Let denote the subtree of composed solely of candidate tokens. The maximum depth of is .
Proposition B.1.
For a sparse tree state , where each candidate token follows a path from the root, and the acceptance probability at each path position , the expected number of tokens generated is given by , where represents the contribution of a token to the expected number of tokens.
We then propose an approximation of the amortized number of tokens generated, by considering the tokens generated at the current and the next decoding step.
Proposition B.2.
The expected total number of tokens generated for the sparse tree state at the current and the next decoding step is given by , where represents the state transition probability from state to state .
We are now ready to introduce Proposition B.3, which we use in the pruning algorithm.
Proposition B.3.
For a sparse tree state with candidate subtree , the change in expected total tokens due to the removal of a prompt token at candidate token is given by , where is the acceptance probability of candidate , denotes the number of prompt tokens prior to removal. We assume that .
We now introduce the formulation of the real amortized number of tokens generated.
Proposition B.4.
The amortized number of tokens generated for the sparse tree state is given by , where is the steady-state probability of state , and is the function defined in Proposition B.1.
The sparse tree construction algorithm can now be formulated as finding the sparse tree with candidate tokens and prompt tokens to maximize :
For a fixed tree size , we explore all combinations of and where , to identify the sparse tree that maximizes .
Appendix C Comparison with More Baselines
C.1 Comparison with Speculative Decoding
| Temperature | Method | Speedup Ratio |
| T=1 (Stochastic) | PPD | 2.26 |
| Eagle | 2.15 | |
| Sps | 1.38 | |
| T=0 (Greedy) | PPD | 2.15 |
| Eagle | 2.60 | |
| Sps | 1.68 |
Table 3 compares the speedup ratios of different speculative decoding methods on Vicuna-7B and MT-Bench. PPD outperforms Eagle at , whereas Eagle achieves better performance at . This trend aligns with findings in typical acceptance Cai et al. (2024), which tends to be more effective at higher temperatures.
PPD serves as a cost-effective and memory-efficient alternative to speculative decoding Li et al. (2024b); Kim et al. (2024), specifically designed for edge and mobile deployments. Both memory usage and training time are critical metrics in this context. More free memory enables the use of larger batch sizes, resulting in improved throughput.
Memory efficiency is also vital for long-context tasks, where a large KV cache size could otherwise lead to OOM errors. Additionally, reducing training time significantly lowers training costs, making PPD more accessible for local deployments. Faster training is particularly valuable in online learning scenarios, where the target model distribution may shift over time, requiring frequent retraining.
C.2 Comparison with More Baselines
| Temperature | Method | Speedup Ratio |
| T = 0.1 | PPD | 2.28 |
| PaSS | 1.47 | |
| T = 0.8 | PPD | 2.40 |
| PaSS | 1.24 |
| Method | Speedup Ratio |
| PPD | 2.26 |
| Draft&Verify | 1.456 |
Appendix D Training Loss
We study the training loss of PPD with different EPTs. Figure 9(a) shows that, with 3 prompt tokens and 1 EPT, the initial loss is quite high, starting above 5. There is a sharp decrease in loss within the first epoch, dropping below 2. After this initial drop, the loss stabilizes and oscillates around a value slightly below 2 for the remainder of the training epochs (up to epoch 12). The loss oscillations remain within a narrow range, indicating consistent performance. The fluctuation can be attributed to the insertion of prompt tokens at random positions. On the other hand, Figure 9(b), with 3 prompt tokens and 100 EPTs, shows the initial loss starting below 3, significantly lower than PPD with 1 EPT. Similarly, there is a sharp decrease within the first epoch, with the loss dropping to around 2.5. However, unlike PPD with 1 EPT, the loss continues to decrease gradually over the epochs, showing a downward trend. This suggests that increasing the number of EPTs improves the model’s learning capacity and reduce training loss more effectively over time.
Appendix E Generalizability of Prompt Tokens to Different Tasks
While original prompt tuning tailors LLMs for specific downstream tasks, our prompt tokens are task-agnostic. To demonstrate their generalizability, Figure 10 shows the prediction accuracy of a single set of prompt tokens across three different datasets. Trained on the ShareGPT dataset, these tokens generalize effectively to unseen tasks. We also report the prediction accuracy for using 5 prompt tokens.
Appendix F Extended Ablation Study
F.1 Effect of EPTs on Prediction Accuracy
| EPT | @1 Top-1 | @1 Top-5 | @2 Top-1 | @2 Top-5 |
| 100 | 0.506 | 0.794 | 0.276 | 0.602 |
| 50 | 0.502 | 0.791 | 0.281 | 0.604 |
| 20 | 0.501 | 0.791 | 0.276 | 0.607 |
| 10 | 0.494 | 0.786 | 0.273 | 0.600 |
| 5 | 0.499 | 0.787 | 0.265 | 0.596 |
| 2 | 0.486 | 0.777 | 0.259 | 0.583 |
| 1 | 0.472 | 0.771 | 0.248 | 0.576 |
Table 6 presents the prediction accuracy of PPD using different EPTs. The results indicate that increasing the number of EPTs generally enhances the prediction accuracy of PPD, particularly for long-range token predictions. Higher EPT numbers (e.g., 100 and 50) consistently produce better prediction accuracy compared to lower EPT numbers.
F.2 Impact of Knowledge Distillation (KD), Epochs, and Batch Size on Prediction Accuracy
| EPT | KD | Epoch | Batch | @1 Top-1 | @1 Top-5 | @2 Top-1 | @2 Top-5 |
| 100 | Yes | 1 | 4 | 0.504 | 0.793 | 0.273 | 0.598 |
| 100 | Yes | 2 | 4 | 0.512 | 0.797 | 0.288 | 0.611 |
| 100 | Yes | 6 | 4 | 0.520 | 0.802 | 0.302 | 0.620 |
| 100 | Yes | 8 | 4 | 0.524 | 0.804 | 0.307 | 0.619 |
| 100 | Yes | 10 | 4 | 0.523 | 0.804 | 0.305 | 0.623 |
| 100 | Yes | 12 | 4 | 0.525 | 0.805 | 0.308 | 0.625 |
| 100 | No | 12 | 4 | 0.506 | 0.794 | 0.276 | 0.602 |
| 100 | Yes | 12 | 1 | 0.530 | 0.809 | 0.309 | 0.626 |
| 1 | Yes | 12 | 1 | 0.484 | 0.775 | 0.259 | 0.581 |
| 1 | Yes | 2 | 4 | 0.474 | 0.773 | 0.247 | 0.574 |
| 1 | Yes | 6 | 4 | 0.480 | 0.773 | 0.250 | 0.580 |
| 1 | Yes | 8 | 4 | 0.484 | 0.778 | 0.257 | 0.583 |
| 1 | Yes | 10 | 4 | 0.482 | 0.777 | 0.257 | 0.584 |
| 1 | Yes | 12 | 4 | 0.485 | 0.779 | 0.261 | 0.586 |
| 1 | No | 12 | 4 | 0.472 | 0.771 | 0.248 | 0.576 |
Table 7 summarizes our results with different settings. We analyze the effect of each factor on the prediction accuracy in the following discussion.
F.2.1 Training Epochs
We first investigate the effect of the number of training epochs on prediction accuracy. For models using 100 EPTs with KD enabled and a batch size of 4, we observe a steady improvement in prediction accuracy as the number of epochs increases. Specifically, the Top-1 accuracy at a 1-token distance increases from 0.504 at 1 epoch to 0.525 at 12 epochs, while the Top-5 accuracy at a 1-token distance improves from 0.793 to 0.805. Similarly, Top-1 accuracy at a 2-token distance increases from 0.273 to 0.308, and Top-5 accuracy at a 2-token distance improves from 0.598 to 0.625 over the same range of epochs. This trend demonstrates the positive impact of prolonged training on the performance of PPD when KD is applied.
F.2.2 Knowledge Distillation
When KD is not applied, as shown for 100 EPTs at 12 epochs with a batch size of 4, the performance metrics are generally lower. The improvement in prediction accuracy with KD is up to 12%. This suggests that KD contributes significantly to prediction accuracy for PPD.
F.2.3 Effect of Batch Size
We also examine the impact of batch size on the prediction accuracy. For the model trained with 100 EPTs, KD enabled, and 12 epochs, reducing the batch size from 4 to 1 results in a slight improvement in prediction accuracy up to 1%.
F.3 Prefix Tuning + Prompt Token
Prefix tuning (Li and Liang, 2021), similar to prompt tuning, provides a parameter-efficient approach to fine-tune a pre-trained model. Unlike prompt tuning, it modifies the KV cache of every attention layer by prepending trained vectors. We hypothesize that the combination of prefix tuning and prompt tokens can lead to greater learning capacity and higher prediction accuracy. This hypothesis is based on the intuition that prompt tokens should see a different context than the input tokens when predicting long-range tokens. For example, if the input sequence is "Once upon a time", then enhancing the input with a prompt template might provide more suitable semantic context for long-range prediction. An enhanced input like "Predict the next-next token. Once upon a time" might empower the prompt token to predict the correct next-next token. Prefix tuning serves as the prompt template to enhance the hidden states visible to the prompt tokens.
To retain the original model’s distribution, we modify the attention mask so that prefix tokens are only visible to prompt tokens. This ensures that we can generate outputs that preserve the original model’s distribution. We posit that prompt tokens at different positions should see different contexts so we allow a prompt token at a specific position to see a distinct set of prefix tokens, as shown in Figure 11.
| Prefix Tuning | @1 Top-1 | @1 Top-5 | @2 Top-1 | @2 Top-5 |
| No | 0.485 | 0.779 | 0.261 | 0.586 |
| Yes | 0.412 | 0.738 | 0.204 | 0.541 |
Table 8 compares the prediction accuracy of PPD with and without the use of prefix tuning. The results show that the models without prefix tuning outperform those with prefix tuning up to 28%, which suggests that, in this setup, prefix tuning does not enhance the prediction accuracy of PPD. Instead, it appears to degrade performance, potentially due to the complexity introduced by modifying the KV cache of attention layers with the prefix token. Unlike prompt tokens, prefix tokens do not interact with input tokens, meaning they do not change dynamically through the transformer layers based on the input context. This lack of interaction and dynamic adjustment could be a factor contributing to the decreased prediction accuracy observed with prefix tuning.
F.4 Custom Decoding Heads + Prompt Token
It has been demonstrated that a fine-tuned decoding head alone can effectively predict long-range tokens (Stern et al., 2018; Cai et al., 2024). Thus, we hypothesize that combining a separately fine-tuned decoding head with prompt tokens might further enhance the potential of PPD. As shown in Figure 12, we trained a separate decoding head to transform only the hidden states of prompt tokens into logits. A key distinction from Medusa is that this decoding head is responsible for generating tokens at multiple positions, rather than just one.
We propose two training methods. In the first method, the custom decoding head and prompt tokens are trained together from scratch in a single stage. In the second method, the prompt tokens are initially trained for 2 epochs, followed by training both the prompt tokens and the decoding head with a smaller learning rate in a two-stage process.
| Method Name | @1 Top-1 | @1 Top-5 | @2 Top-1 | @2 Top-5 |
| PPD without custom decoding head | 0.485 | 0.779 | 0.261 | 0.586 |
| PPD with custom decoding head (1-stage) | 0.385 | 0.614 | 0.229 | 0.482 |
| PPD with custom decoding head (2-stage) | 0.506 | 0.795 | 0.276 | 0.602 |
Table 9 presents the prediction accuracy of PPD with and without a custom decoding head. When trained using the single-stage method, PPD with the custom decoding head shows a 12%-21% decrease in prediction accuracy compared to the baseline PPD without the custom decoding head. This suggests that the single-stage approach does not result in stable or effective training.
In contrast, the two-stage training method results in a limited improvement of 2.1%-4.3% in prediction accuracy compared to the baseline. This suggests that adding a custom decoding head may not be necessary, given the additional trainable parameters and the limited improvement in prediction accuracy.
F.5 Attention Masking for EPTs
In this paper, we proposed a specialized attention mask for EPTs to achieve the effect of prompt ensemble. However, there are alternative masking strategies available. Here, we describe and compare three types of attention masks that we implemented and experimented with.
F.5.1 Ensemble Attention Masking
The ensemble attention masking is the masking strategy we previously described. In this approach, EPTs are divided into disjoint groups, where is the number of EPTs per prompt token. All EPTs across prompt tokens are placed in the same group. An EPT in group can only attend to EPTs that meet the following two criteria: 1) they must belong to group , and 2) their position indices must be smaller than the position index of . Since this masking strategy effectively averages the results of disjoint groups of EPTs, we refer to it as the "ensemble attention masking". Figure 13(a) provides an example of the ensemble attention masking.
F.5.2 Decoder-like Attention Masking
Decoder-like attention masking is a simple strategy where EPTs can only attend to EPTs with smaller position indices. This results in a triangular-shaped attention mask, similar to the one used in decoder layers, hence the name "decoder-like attention masking". Figure 13(b) provides an example of this masking strategy.
F.5.3 Encoder-like Attention Masking
In encoder-like attention masking, an EPT corresponding to a prompt token can attend to all EPTs with smaller position indices as well as all EPTs associated with . This allows EPTs to see both preceding and succeeding EPTs, similar to the token visibility in an encoder layer, hence the name "encoder-like attention masking". Figure 13(c) illustrates this masking strategy.
F.5.4 Results
| Method Name | @1 Top-1 | @1 Top-5 | @2 Top-1 | @2 Top-5 |
| PPD with ensemble attention mask | 0.506 | 0.794 | 0.276 | 0.602 |
| PPD with decoder attention mask | 0.465 | 0.755 | 0.262 | 0.572 |
| PPD with encoder attention mask | 0.473 | 0.765 | 0.256 | 0.573 |
The results in Table 10 indicate that the ensemble attention mask outperforms the other masking strategies. In comparison, the PPD with decoder attention mask shows 4.9%-8.0% lower prediction accuracy. The PPD with encoder attention mask also underperforms in prediction accuracy relative to the ensemble attention mask by 3.7%-7.2%.
These results suggest that the ensemble attention mask is the most effective strategy among the three, likely due to its ability to effectively average the votes of disjoint groups of EPTs, thereby improving prediction accuracy. The decoder-like and encoder-like attention masks, while simpler, do not provide the same level of performance, indicating that the structure and specificity of the ensemble attention mask better facilitate accurate long-range token prediction. Additionally, ensemble attention masking is more sparse, which offers greater potential for optimization.
F.6 Aggregation Method for EPTs
In addition to simply averaging the logits from EPTs, we explored more advanced aggregation methods. For instance, we applied learned weights to aggregate the logits. The final logit can be expressed as:
where is the number of EPTs and is the learned scalar weight for the EPT.
| Aggregation Method | @1 Top-1 | @1 Top-5 | @2 Top-1 | @2 Top-5 |
| Average | 0.506 | 0.794 | 0.276 | 0.602 |
| Learned Weight | 0.503 | 0.779 | 0.250 | 0.576 |
The results in Table 11 show the prediction accuracy of PPD with two different aggregation methods for EPTs: simple averaging and learned weights. When using learned weights to aggregate logits, the model shows a slight decrease of 0.6%-9.4% in prediction accuracy.
These results suggest that while learned weights provide a more flexible aggregation method, they do not necessarily lead to improved prediction accuracy in this context. The simplicity and stability of the averaging method appear to offer better performance, possibly due to the additional complexity and potential overfitting introduced by learning the weights.
F.7 Multi-exit Ensemble
While using EPTs for prompt ensemble improves prediction accuracy, it also increases input length, resulting in higher computational overhead and forward pass latency. To address this, we propose the use of a multi-exit ensemble method. In multi-exit ensemble, the hidden states of a prompt token from the last decoder layers are extracted and averaged to produce the final hidden state, which is then decoded by the decoding head into a guess token, as illustrated in Figure 14. This approach achieves prompt ensemble without the associated computational costs.
The hypothesis is that taking the hidden states from the last few decoder layers for ensemble might work because these layers capture increasingly abstract and high-level representations of the input sequence. By averaging the hidden states from multiple layers, we can combine diverse but complementary information, leading to a more robust and accurate final hidden state. Additionally, since the final layers are closest to the output, they are more likely to contain refined and contextually relevant information, making the ensemble more effective.
| Method Name | @1 Top-1 | @1 Top-5 | @2 Top-1 | @2 Top-5 |
| PPD without multi-exit | 0.485 | 0.779 | 0.261 | 0.586 |
| PPD with 3 exits | 0.422 | 0.723 | 0.214 | 0.517 |
| PPD with 2 exits | 0.420 | 0.723 | 0.213 | 0.518 |
Table 12 shows the comparison of prediction accuracy of PPD with and without mult-exit ensemble. The results indicate that the introduction of multi-exit ensemble with both 2 and 3 exits results in a 7%-18% decrease in prediction accuracy compared to the baseline model without multi-exit.
These findings suggest that the multi-exit ensemble approach, as implemented, does not enhance prediction accuracy and instead leads to a notable decrease in performance. This may be due to the averaging of hidden states from multiple layers introducing noise or reducing the specificity of the representations needed for accurate prediction. Further refinement of the multi-exit ensemble may be necessary to achieve the desired improvements in accuracy.
Appendix G Effect of Batch Size on Speedup
| Batch Size | 1 | 2 | 3 | 4 |
| PPD Speedup Ratio (w/o Tree Attention) | 1.71 | 1.65 | 1.63 | 1.64 |
| PPD Speedup Ratio (with Tree Attention) | 2.26 | 1.90 | 1.58 | 1.52 |
As shown in Table 13, consistent speedup ratios are achieved across different batch sizes without tree attention. However, with tree attention, the speedup ratio decreases as batch size increases, a pattern similar to other parallel and speculative decoding methods.
Appendix H Experiment Details
For the throughput experiments, each result is obtained by averaging three separate runs. The standard deviations of these runs are reported as error bars in the bar charts. To ensure a fair comparison in our comparative experiments, we maintained consistent hardware settings and software versions.
We selected 3 prompt tokens because adding more would not further increase the expected acceptance length due to the tree size limit. The number of EPTs per prompt token was optimized to maximize throughput.
In Fig. 2, the temperature settings for PPD, Eagle (Li et al., 2024b), and Medusa (Cai et al., 2024) follow the default configuration, while the other models use a greedy setting (temperature=0). This choice is based on findings that retrieval-based methods perform significantly worse in non-greedy settings. Similarly, LOOKAHEAD DECODING (Fu et al., 2024), REST (He et al., 2023), and PLD (Saxena, 2023) in Fig. 4 also use a temperature setting of 0 for the same reasons.
Appendix I Societal Impact
In this paper, we proposed PPD to accelerate LLMs easily and cheaply. Since PPD reduces the time required for handling a single inference request, it could bring down the cost of deploying LLMs for both the companies and the public. This might lead to increased accessibility of LLM services. Moreover, latency-sensitive applications like chatbots will benefit greatly from the usage of PPD as it reduces the inference latency greatly, thereby enhancing the user experience.
While PPD aims to make AI more accessible, there may still be a digital divide where certain communities lack the necessary infrastructure, such as stable internet connections or modern hardware, to fully benefit from these advancements. This could further widen the gap between technology-privileged and underserved populations. On the other hand, PPD might be misused by malicious parties to manipulate the output of the original LLM, resulting in the generation of unreliable information and fake data.